OpenStudy (anonymous):

Six divided by the difference of a number and 2 minus 5 divided by a number plus 2, equals 3 times the reciprocal of the difference of the number squared and 4. What is the number?

OpenStudy (anonymous):

@AJ01 @Reptile_Girl777 @goatgal102 @Godlovesme @E>Nattie12345 @Quan99 @brittney_abshire @kittycat123 @vera_ewing @farmgirl11 @xo.minnie.xox @MakaylaTracy

OpenStudy (anonymous):

6/(x-2)

OpenStudy (anonymous):

5/(x+2)

OpenStudy (anonymous):

3/(x^2-4)

OpenStudy (anonymous):

6/(x-2) +5/(x+2)=3/(x^2-4)

OpenStudy (anonymous):

6/(x-2)+5/(x+2)=3/((x-2)(x+2))

OpenStudy (anonymous):

multiply by both side (x^2-4)

OpenStudy (anonymous):

6(x+2)+5(x-2)=3

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

1/11? Thank you so much!